OpenStudy (anonymous):

a/4-21=7 i got 112 i just wana make sure i did it right

OpenStudy (anonymous):

Evaluate a=112 112/4-21=7 if it's true you did right. Use the calculator if you don't want to do the math handwritting.

OpenStudy (anonymous):

yes a=112

OpenStudy (anonymous):

typ some thing so i can give you a medal lol

OpenStudy (anonymous):

so it would be 28-21=7

OpenStudy (anonymous):

yes 28-21=7

OpenStudy (anonymous):

Okk thank you everyone :P

OpenStudy (anonymous):

can i have a medal too =)

OpenStudy (anonymous):

and do you need more help

OpenStudy (anonymous):

thankx =)